# Made-to-Measure Outperforms Schwarzschild’s Method

### v1

@@ Abstract
Syer and Tremaine’s made-to-measure method and Schwarzschild’s orbit superposition method are well-known within the field of stellar dynamical modeling. This research is concerned with assessing and comparing the operational capabilities of the two methods and, in particular, the impact on observable reproduction, orbit classifications and computer elapsed times when using low orbit numbers (8000 orbits) with different observational data sets and initial conditions. Both methods are able to reproduce observed data with mean $\chi^{2}\approx 1$ or less. However, the made-to-measure process does so three to five times faster than the orbit superposition method, and this starts to make the made-to-measure process attractive for analysing galaxy surveys. For a given set of initial conditions, both methods produce similar orbit classifications but the orbits behind the classifications are not the same. Orbits which are common do not have the same weights. Different initial conditions result in different classifications.
@@ 1 Introduction
Syer and Tremaine’s made-to-measure method (Syer1996) (M2M) and Schwarzschild’s orbit superposition method (Schwarz1979) (SCHW) are well-known within the field of stellar dynamical modeling. Of the two, SCHW has more citations (885) at the time of preparing this article (mid-2025) than M2M (168). M2M and SCHW have been compared previously in Long2012 where M2M and SCHW were shown to be able to produce similar estimates of galaxy mass-to-light ratios. It is now appropriate that a further comparison is made.
@@
This article will not address the third factor: this may be better left to machine learning. The first two factors will be considered further. At the time of the first comparison, the accepted assessment was that M2M, with its ability to use significantly more orbits (numbers of $10^{5}$ to $10^{6}$ were not uncommon), must of course be producing the better models. SCHW by comparison was claimed to model real galaxies with only a few hundred orbits. The question being posed in this article is how does M2M compare with SCHW when run using fewer orbits, for example $\approx 10^{4}$. The objective of this research therefore is to assess and compare the operational capabilities of M2M and SCHW, in particular the impact on observable reproduction, orbit classifications and computer elapsed times, when using low orbit numbers.
@@ 2 Approach
More information is provided in Sections 3 and 4.
It is not intended that the behavior and performance of all the capabilities of M2M will be compared with all the capabilities of SCHW but only the core capabilities concerned with luminosity and kinematical data reproduction. Features such as discrete data handling, or chemical modeling, or orbit classification constraints are out of scope, as is parameter estimation using the methods. Note also that only axisymmetric (oblate) models are utilized: the investigations are not affected by choice of gravitational potential. The kinematical observables are varied so that both velocity moments and Gauss-Hermite coefficients are modeled - see Section 3. The earlier comparison in Long2012 used Gauss-Hermite coefficients only.
SCHW requires the use of NNLS software to create orbit weights, We choose NNLS software that implements the Lawson and Hanson active set method (LH1974) since it appears from the literature to be the most used NNLS implementation for SCHW. The active set algorithm deliberately zeroises the weights of any orbits it chooses not to use in reproducing the observed data, and this is taken into account in the analysis software used.
Model selection frameworks such as Lipka2021 and Thomas2022 are not used. Some use is made of an experimental Gaussian processes approach to M2M hyper-parameter determination.
Other than to compare M2M and SCHW orbit classifications for IllustrisTNG data with Xu2019, no external comparisons of models are made.
@@ 3 Observational Data
The observational data sets used in this article are the same as those used in Long2025. The data tables here are modified versions of the tables from that paper.
For the ATLAS3D data set (AtlasI), data from four galaxies are utilized - NGC 1248, NGC 3838, NGC 4452, and NGC 4551. These galaxies have various inclinations, different mass-to-light ratios, and sense of bulk rotation (see Table 2). Surface brightness values are calculated using multi-Gaussian expansions (Sect. 4.1). The IFU cells for kinematical data (first and second velocity moments) are the Voronoi cells resulting from signal-to-noise processing of the raw data.
@@
The inclinations are taken from AtlasXV, while the M/L ratios were determined by M2M modeling in Long2016. The SB max radius column gives the overall size of the polar grid used for surface brightness data values. The IFU column gives the number of Voronoi cells for the kinematical data.
For the IllustrisTNG galaxies (Nelson2019), five axisymmetric galaxies (A0490, A1090, A1190, A1290, and A1390) are utilized, and are viewed edge on with the mass-to-light ratio taken as $1.0$ for all galaxies. Surface brightness values are again calculated using multi-Gaussian expansions. The IFU cells for kinematical data (the Gauss-Hermite coefficients $h_{1}$ to $h_{4}$) are constructed to represent a real galaxy’s data (see Table 3).
@@ 4 Methods
With the exception of M2M, all the methods and techniques in this section are very similar to those in Long2025 and are not described in detail here. The M2M implementation was used recently in Long2021, and is described more fully in Long2016 and other papers by the lead author prior to that.
@@ 4.1 Gravitational Potentials
The gravitational potentials and their derivatives used for modeling are derived from MGEs of the surface brightness data (Emsellem1994; Cappellari2002). For the IllustrisTNG simulated data, dark matter potentials are included as spherical Gaussians, and no black hole modeling takes place. For the ATLAS3D data, no dark matter is included for consistency with AtlasXV, and central black holes are modeled as point sources.
@@ 4.3 Orbit Classifications
The circularity measure $\lambda_{z}$ (Zhu2018) is used to classify orbits into one of hot ($\left|\lambda_{z}\right|<=0.25$), cold ($\lambda_{z}>=0.80$), warm ($0.25<\lambda_{z}<0.8$) or counter-rotating. The total orbit weights in each of the classes are taken as the overall orbit classification for a galaxy. Comparing orbit classifications produced by M2M and SCHW is a key part of our modeling analysis. Note that orbit classifications are not used as constraints themselves, as in Long2025. However, the sum of the weights is constrained to be equal to one, and so the sum of the orbit classes must also equal one.
@@ 4.7 Software Utilization
As in Long2025, the software base for constructing initial conditions and for M2M and SCHW modeling is the lead author’s implementation of the Syer1996 M2M stellar dynamical modeling method together with the Schwarz1979 orbit superposition method (a single code set supports both methods). This software was first used in Long2010, and, as indicated, more recently in Long2025. The software is Python 3 based with some use of Cython (behnel2010cython) and C for performance critical code. The MPI message passing interface is used to achieve parallelization of orbit integration, with the same number of cores being used for both M2M and SCHW. The optimization software used for NNLS is taken from the SCIPY package version 1.16 (2020SciPy-NMeth) (earlier versions may have NNLS performance issues). The Gaussian processes application is based on routines within the publicly available scikit-optimize package. Using a $20$ core desktop PC, multi-processor working is employed with up to 10 cores active in a modeling run.
@@ 5.2 $\chi^{2}$ and Elapsed Time Comparisons between M2M and SCHW
The focus in this section is on how well observed data are reproduced, and comparing modeling elapsed times. Orbit classifications are considered in the next section. M2M and SCHW models using both 3I and MDJV initial conditions and both data sets are created. Table 4 contains the results for the ATLAS3D galaxy data, and Table 5 for the IllustrisTNG galaxy data. In the tables, all the mean observable $\chi^{2}$ values are $\approx 1$ or less indicating that the observed data have been well-reproduced by both M2M and SCHW. An example is shown in Appendix A. M2M weight convergence values are $93\%$ implying that convergence to constant weights, as expected in Syer1996, is close to being achieved. For ATLAS3D , the modeling elapsed times show that M2M performs faster than SCHW by a factor of $\approx 4$ for 8000 orbit models and by $\approx 5$ for 16000 orbit models. For IllustrisTNG data, M2M outperforms SCHW by a factor of $\approx 3$. Any of these performance improvement factors would benefit large surveys where there are thousands of galaxy data sets to be processed. The NNLS contributions to SCHW elapsed times show that, with increasing model size (numbers of orbits in this case), the performance difference will become more marked depending on the algorithms within the NNLS implementation.
@@ 5.3 Orbit Classification Comparisons between M2M and SCHW
The comparisons between orbit classification results from modeling with M2M and SCHW are shown in Table 6. M2M and SCHW are producing similar orbit classifications from the same observed data and initial conditions, with ATLAS3D and MDJV showing the most variation in classification and IllustrisTNG and MDJV showing the least.
@@
M2M and SCHW common orbits weight comparisons for ATLAS3D data (Section 5.3). 3I initial conditions are used for the models. The left hand column shows scatter plots of the weights of the orbits, with the blue dashed line indicating equality of the orbit weights. The key point to note is that the weights are not tightly positioned close to the equality line: even though the orbits used are the same, M2M and SCHW are in general not allocating the same weights to the orbits. The Bland-Altman plots (BA1983) of the average log weight for an orbit against the log difference between its M2M weight and its SCHW weight show that the mean log difference (red solid line) is approximately zero indicating that there is no significant bias towards either method. In addition, weight difference is larger at lower average weights than with larger average weights. The red dashed lines show the $2.5\%$ and $97.5\%$ percentiles with $95\%$ of the difference ratios lying between the lines. The right hand plots show the distributions of the weight ratios. The red lines have the same meaning as in the Bland-Altman plots.
@@ 5.4 Comparisons between 3I and MDJV
5.5 Comparison with Xu2019
The orbit classification comparisons for M2M and SCHW using IllustrisTNG data in Section 5.3 are compared with Xu2019, and the results recorded in Table 9. Apart from the A0490 M2M model, all the other models have differences with Xu2019 in the hot, warm and cold classifications and this is considered further in the Discussion (Section 6). Apart from the counter-rotating classification for the A0490 SCHW model, all other counter-rotating classifications agree, to within $3\%$, with Xu2019. The differences in the MDJV cold column reflect the fact that MDJV produces fewer cold orbits than 3I, and does not meet the orbital needs of the observed data.
Table 9: IllustrisTNG M2M and SCHW Comparisons with Xu2019 - Orbit Classifications
3I Initial Conditions Xu2019 MDJV Initial Conditions
@@ 5.5 Comparison with Xu2019
Orbit classification comparisons for M2M and SCHW with Xu2019 (see Section 5.5). Comparisons in the table are made horizontally between models and Xu2019. Classification differences of $\geq 3\%$ with Xu2019 are indicated in red and blue. Apart from the A0490 M2M model, all the other models have some differences with Xu2019. The values in the counter-rotating columns are notable for their lack of differences. The MDJV cold column reflects differences between the initial conditions where MDJV has fewer cold orbits than 3I.
@@ 6 Discussion
    Both M2M and SCHW using IllustrisTNG data as in Xu2019 do not produce orbit classifications which match those in Xu2019 (see Table 9). It therefore appears that the gravitational potential, observed data sets and initial conditions used by the methods are not able to produce good ‘orbit level’ representations of the associated galaxies.
@@
All comparisons between M2M and SCHW will, at some level, be implementation specific, but there seems to be little to be gained by seeking to address known modeling shortfalls in an SCHW-like context. M2M is inherently more flexible and scalable than SCHW. Orbit weights are available within the model from the outset so more complex observables are easier to model or estimate. Data uncertainties are an integral part of the M2M weight determination process, and are accommodated far more so than SCHW does in reaching the final solution. Once designed, achieving a basic M2M implementation in approximately six to eight weeks can be accomplished by an experienced programmer using a high level language such as Python, and including use of parallelization. Creating a GPU based implementation for M2M can be achieved using tooling such as PyTorch (paszke2019). However to gain the full performance benefits, this will require careful assessment and design of the GPU parallelizable features of M2M. Given that M2M employs a number of techniques that are used in machine learning, perhaps they might provide an alternative way forward and this will be investigated in the future by the authors
@@ 7 Conclusions
The objective set out in the Introduction to assess and compare the operational capabilities of M2M and SCHW has been met. Both methods are able to replicate observed data to an acceptable level but M2M outperforms SCHW being several times faster. This speed improvement makes M2M more desirable in a galaxy survey context than SCHW. Both methods produce similar orbit classifications for a given set of initial conditions, but the classifications do appear to be particular to the initial conditions used. In addition, the weights of the common orbits under-pinning the classifications differ and are method specific. Looking ahead, M2M is the more attractive option currently for stellar dynamical modeling in the 21st century, but could eventually be overtaken by machine learning based methods. However, in the mean time, this paper acts as a flag to the stellar dynamical modeling community that M2M should be seriously considered as a way of alleviating some of the shortcomings of SCHW.

### v2

@@ Abstract
Syer and Tremaine’s made-to-measure method and Schwarzschild’s orbit superposition method are well-known within the field of stellar dynamical modeling. This research is concerned with assessing and comparing the operational capabilities of the two methods and, in particular, the impact on observable reproduction, orbit classifications and computer elapsed times when using low orbit numbers (8000 orbits) with different observational data sets and initial conditions. Both methods are able to reproduce observed data with mean $\chi^{2}\approx 1$ or less. However, the made-to-measure process does so three to five times faster than the orbit superposition method, and this starts to make the made-to-measure process attractive for analyzing galaxy surveys. For a given set of initial conditions, both methods produce similar orbit classifications but the orbits behind the classifications are not the same. Orbits which are common do not have the same weights. Different initial conditions result in different classifications.
@@ 1 Introduction
Syer and Tremaine’s made-to-measure method (Syer and Tremaine, 1996) (M2M) and Schwarzschild’s orbit superposition method (Schwarzschild, 1979) (SCHW) are well-known within the field of stellar dynamical modeling. Of the two, SCHW has more citations (885) at the time of preparing this article (mid-2025) than M2M (168). M2M and SCHW have been compared previously in Long and Mao (2012) where M2M and SCHW were shown to be able to produce similar estimates of galaxy mass-to-light ratios. It is now appropriate that a further comparison is made.
@@
This article will not address the third factor: this may be better left to machine learning. The first two factors will be considered further. At the time of the first comparison, the accepted assessment was that M2M, with its ability to use significantly more orbits (numbers of $10^{5}$ to $10^{6}$ were not uncommon), must of course be producing the better models. SCHW by comparison was claimed to model real galaxies with only a few hundred orbits, and as a consequence was perceived to be much faster than M2M. M2M, however, with the orbit weights being available from the start of modeling is inherently more flexible than SCHW. The question being posed in this article is how does M2M compare with SCHW when run using fewer orbits, for example $\approx 10^{4}$. The objective of this research therefore is to assess and compare the operational capabilities of M2M and SCHW, in particular the impact on observable reproduction, orbit classifications and computer elapsed times, when using low orbit numbers. This is believed to be the first time that such a comparison between the two methods, under controlled conditions, has been undertaken.
@@ 2 Approach
More information is provided in Sections 3 and 4. Overall, the comparison has been designed to be between two functionally equivalent methods with implementation and feature usage differences minimized.
It is not intended that the behavior and performance of all the capabilities of M2M will be compared with all the capabilities of SCHW but only the core capabilities concerned with luminosity and kinematical data reproduction. Features such as discrete data handling, or chemical modeling, or orbit classification constraints are out of scope, as is parameter estimation using the methods. Note also that only axisymmetric (oblate) models are utilized: the investigations are not affected by choice of gravitational potential. The kinematical observables are varied so that both velocity moments and Gauss-Hermite coefficients are modeled - see Section 3. The earlier comparison in Long and Mao (2012) used Gauss-Hermite coefficients only.
SCHW requires the use of NNLS software to create orbit weights, We choose NNLS software that implements the Lawson and Hanson active set method (Lawson and Hanson, 1974) since it appears from the literature to be the most used NNLS implementation for SCHW. The active set algorithm deliberately zeroises the weights of any orbits it chooses not to use in reproducing the observed data, and this is taken into account in the analysis software used.
Model selection frameworks such as Lipka and Thomas (2021) and Thomas and Lipka (2022) are not used. Some use is made of an experimental Gaussian processes approach to M2M hyper-parameter determination.
Other than to compare M2M and SCHW orbit classifications for IllustrisTNG data with Xu et al. (2019), no external comparisons of models are made.
@@ 3 Observational Data
The observational data sets used in this article are the same as those used in Long (2025). The data tables here are modified versions of the tables from that paper.
For the ATLAS3D data set (Cappellari et al., 2011), data from four galaxies are utilized - NGC 1248, NGC 3838, NGC 4452, and NGC 4551. These galaxies have various inclinations, different mass-to-light ratios, and sense of bulk rotation (see Table 2). Surface brightness values are calculated using multi-Gaussian expansions (Sect. 4.1). The IFU cells for kinematical data (first and second velocity moments) are the Voronoi cells resulting from signal-to-noise processing of the raw data.
@@
The inclinations are taken from Cappellari et al. (2013), while the M/L ratios were determined by M2M modeling in Long (2016). The SB max radius column gives the overall size of the polar grid used for surface brightness data values. The IFU column gives the number of Voronoi cells for the kinematical data.
For the IllustrisTNG galaxies (Nelson et al., 2019), five axisymmetric galaxies (A0490, A1090, A1190, A1290, and A1390) are utilized, and are viewed edge on with the mass-to-light ratio taken as $1.0$ for all galaxies. Surface brightness values are again calculated using multi-Gaussian expansions. The IFU cells for kinematical data (the Gauss-Hermite coefficients $h_{1}$ to $h_{4}$) are constructed to represent a real galaxy’s data (see Table 3).
@@ 4 Methods
With the exception of M2M, all the methods and techniques in this section are very similar to those in Long (2025) and are not described in detail here. The M2M implementation was used recently in Long et al. (2021), and is described more fully in Long (2016) and other papers by the lead author prior to that.
@@ 4.1 Gravitational Potentials
The gravitational potentials and their derivatives used for modeling are derived from MGEs of the surface brightness data (Emsellem et al., 1994; Cappellari, 2002). For the IllustrisTNG simulated data, dark matter potentials are included as spherical Gaussians, and no black hole modeling takes place. For the ATLAS3D data, no dark matter is included for consistency with Cappellari et al. (2013), and central black holes are modeled as point sources.
@@ 4.3 Orbit Classifications
The circularity measure $\lambda_{z}$ (Zhu et al., 2018) is used to classify orbits into one of hot ($\left|\lambda_{z}\right|<=0.25$), cold ($\lambda_{z}>=0.80$), warm ($0.25<\lambda_{z}<0.8$) or counter-rotating. The total orbit weights in each of the classes are taken as the overall orbit classification for a galaxy. Comparing orbit classifications produced by M2M and SCHW is a key part of our modeling analysis. Note that orbit classifications are not used as constraints themselves, as in Long (2025). However, the sum of the weights is constrained to be equal to one, and so the sum of the orbit classes must also equal one.
@@ 4.7 Software Utilization
As in Long (2025), the software base for constructing initial conditions and for M2M and SCHW modeling is the lead author’s implementation of the Syer and Tremaine (1996) M2M stellar dynamical modeling method together with the Schwarzschild (1979) orbit superposition method (a single code set supports both methods). This software was first used in Long and Mao (2010), and, as indicated, more recently in Long (2025). The software is Python 3 based with some use of Cython (Behnel et al., 2011) and C for performance critical code. The MPI message passing interface is used to achieve parallelization of orbit integration, with the same number of cores being used for both M2M and SCHW. The optimization software used for NNLS is taken from the SCIPY package version 1.16 (Virtanen et al., 2020) (earlier versions may have NNLS performance issues). The Gaussian processes application is based on routines within the publicly available scikit-optimize package. Using a $20$ core desktop PC, multi-processor working is employed with up to 10 cores active in a modeling run.
@@ 5.2 $\chi^{2}$ and Elapsed Time Comparisons between M2M and SCHW
The focus in this section is on how well observed data are reproduced, and comparing modeling elapsed times. Orbit classifications are considered in the next section. M2M and SCHW models using both 3I and MDJV initial conditions and both data sets are created. Table 4 contains the results for the ATLAS3D galaxy data, and Table 5 for the IllustrisTNG galaxy data. In the tables, all the mean observable $\chi^{2}$ values are $\approx 1$ or less indicating that the observed data have been well-reproduced by both M2M and SCHW. An example is shown in Appendix A. M2M weight convergence values are $93\%$ implying that convergence to constant weights, as expected in Syer and Tremaine (1996), is close to being achieved. For ATLAS3D , the modeling elapsed times show that M2M performs faster than SCHW by a factor of $\approx 4$ for 8000 orbit models and by $\approx 5$ for 16000 orbit models. For IllustrisTNG data, M2M outperforms SCHW by a factor of $\approx 3$. Any of these performance improvement factors would benefit large surveys where there are thousands of galaxy data sets to be processed. The NNLS contributions to SCHW elapsed times show that, with increasing model size (numbers of orbits in this case), the performance difference will become more marked depending on the algorithms within the NNLS implementation.
@@ 5.3 Orbit Classification Comparisons between M2M and SCHW
The comparisons between orbit classification results from modeling with M2M and SCHW are shown in Table 6. M2M and SCHW are producing similar orbit classifications from the same observed data and initial conditions, with ATLAS3D and MDJV showing the most variation in classification and IllustrisTNG and MDJV showing the least. For the same initial conditions, it is to be expected that both M2M and SCHW will yield similar orbit classifications as they have essentially the same log likelihood to maximize to reproduce the observed data.
@@
M2M and SCHW common orbits weight comparisons for ATLAS3D data (Section 5.3). 3I initial conditions are used for the models. The left hand column shows scatter plots of the weights of the orbits, with the blue dashed line indicating equality of the orbit weights. The key point to note is that the weights are not tightly positioned close to the equality line: even though the orbits used are the same, M2M and SCHW are in general not allocating the same weights to the orbits. The Bland-Altman plots (Altman and Bland, 1983) of the average log weight for an orbit against the log difference between its M2M weight and its SCHW weight show that the mean log difference (red solid line) is approximately zero indicating that there is no significant bias towards either method. In addition, weight difference is larger at lower average weights than with larger average weights. The red dashed lines show the $2.5\%$ and $97.5\%$ percentiles with $95\%$ of the difference ratios lying between the lines. The right hand plots show the distributions of the weight ratios. The red lines have the same meaning as in the Bland-Altman plots.
@@ 5.4 Comparisons between 3I and MDJV
5.5 Comparison with Xu et al. (2019)
The orbit classification comparisons for M2M and SCHW using IllustrisTNG data in Section 5.3 are compared with Xu et al. (2019), and the results recorded in Table 9. Apart from the A0490 M2M model, all the other models have differences with Xu et al. (2019) in the hot, warm and cold classifications and this is considered further in the Discussion (Section 6). Apart from the counter-rotating classification for the A0490 SCHW model, all other counter-rotating classifications agree, to within $3\%$, with Xu et al. (2019). The differences in the MDJV cold column reflect the fact that MDJV produces fewer cold orbits than 3I, and does not meet the orbital needs of the observed data.
Table 9: IllustrisTNG M2M and SCHW Comparisons with Xu et al. (2019) - Orbit Classifications
3I Initial Conditions Xu et al. (2019) MDJV Initial Conditions
@@
Orbit classification comparisons for M2M and SCHW with Xu et al. (2019) (see Section 5.5). Comparisons in the table are made horizontally between models and Xu et al. (2019). Classification differences of $\geq 3\%$ with Xu et al. (2019) are indicated in red and blue. Apart from the A0490 M2M model, all the other models have some differences with Xu et al. (2019). The values in the counter-rotating columns are notable for their lack of differences. The MDJV cold column reflects differences between the initial conditions where MDJV has fewer cold orbits than 3I.
@@ 6 Discussion
    Both M2M and SCHW using IllustrisTNG data as in Xu et al. (2019) do not produce orbit classifications which match those in Xu et al. (2019) (see Table 9). It therefore appears that the gravitational potential, observed data sets and initial conditions used by the methods are not able to produce good ‘orbit level’ representations of the associated galaxies. However, the classification values in Xu et al. (2019) are calculated from six dimensional phase space data, and are taken as the true classifications for the simulated galaxies. The observed data for modeling taken from the simulated galaxies do not contain sufficient information to recreate these true classifications. This is consistent with the data from real galaxies. It is possible to use the Xu et al. (2019) classifications as model constraints and this was demonstrated in Long (2025).
@@
All comparisons between M2M and SCHW will, at some level, be implementation specific, but there seems to be little to be gained by seeking to address known modeling shortfalls in an SCHW-like context. M2M is inherently more flexible and scalable than SCHW. Orbit weights are available within the model from the outset so more complex observables are easier to model or estimate. Data uncertainties are an integral part of the M2M weight determination process, and are accommodated far more so than SCHW does in reaching the final solution. Once designed, achieving a basic M2M implementation in approximately six to eight weeks can be accomplished by an experienced programmer using a high level language such as Python, and including use of parallelization. Creating a GPU based implementation for M2M can be achieved using tooling such as PyTorch (Paszke et al., 2019). However to gain the full performance benefits, this will require careful assessment and design of the GPU parallelizable features of M2M. Given that M2M employs a number of techniques that are used in machine learning, perhaps they might provide an alternative way forward and this will be investigated in the future by the authors
@@ 7 Conclusions
The objective set out in the Introduction to assess and compare the operational capabilities of M2M and SCHW has been met. Using the low orbit numbers typical of SCHW, both methods are able to replicate observed data to an acceptable level but M2M outperforms SCHW being several times faster. This speed improvement makes M2M more desirable in a galaxy survey context than SCHW. Both methods produce similar orbit classifications for a given set of initial conditions, but the classifications do appear to be particular to the initial conditions used. In addition, the weights of the common orbits under-pinning the classifications differ and are method specific. Looking ahead, based on the results here and given its inherent flexibility, M2M is the more attractive option currently for stellar dynamical modeling in the 21st century, but could eventually be overtaken by machine learning based methods. In the mean time, this paper acts as a flag to the stellar dynamical modeling community that M2M should be seriously considered as a way of alleviating some of the shortcomings of SCHW.